OpenStudy (anonymous):

WILL FAN AND MEDAL Please help!.!. been trying to work these out for 3 hours.. 2/3*x+4=3/5*x-2 solve for x 3(h-4)=-1/2(24-6h solve for h ax+bc=-c idek what this is.?.?

OpenStudy (phi):

\[ \frac{2}{3} x + 4= \frac{3}{5}x -2 \] one way is, as a first step, to multiply both sides (and *all* terms on both sides) by 15 can you do that?

OpenStudy (anonymous):

15 because its the LCD?right?

OpenStudy (phi):

yes. this technique is called "clearing the denominator" (the other way is to combine fractions)

OpenStudy (anonymous):

ok so 10x+60=9x-30

OpenStudy (phi):

yes, now add -9x to both sides add -60 to both sides simplify

OpenStudy (phi):

10x-9x +60=9x-9x -30 on the right side 9x-9x is 0 on the left side 10 x's take away 9 x's leaves 1x 1x + 60 = 0-30 x + 60 = -30 now add -60 to both sides.

OpenStudy (anonymous):

x=-90?

OpenStudy (anonymous):

the other answers are 9: identity property 10: x=-c/a+b

OpenStudy (phi):

for the last one \[ ax+bc=-c \\ ax + bc - bc = -c - bc \\ ax = -c -bc \] divide both sides by a \[ x = \frac{-c-bc}{a} \] we could factor out -c from the top, and write this as \[ x = \frac{-c(1+b)}{a} \] or in other ways, e.g. \[ x = -(b+1) \cdot \frac{c}{a} \]